# Semantic VLM Dataset for Safe Autonomous Driving

### v1

@@ Abstract
CAR-Scenes is a frame-level dataset for autonomous driving that enables the training and evaluation of vision–language models (VLMs) for interpretable, scene-level understanding. We annotate 5,192 images drawn from Argoverse1, Cityscapes, KITTI, and nuScenes using a 28-key category/sub-category knowledge base covering environment, road geometry, background-vehicle behavior, ego-vehicle behavior, vulnerable road users, sensor states, and a discrete severity scale (1–10), totaling 350+ leaf attributes. Labels are produced by a GPT-4o–assisted vision–language pipeline with human-in-the-loop verification; we release the exact prompts, post-processing rules, and per-field baseline model performance. CAR-Scenes also provides attribute co-occurrence graphs and JSONL records that support semantic retrieval, dataset triage, and risk-aware scenario mining across sources. To calibrate task difficulty, we include reproducible, non-benchmark baselines—notably a LoRA-tuned Qwen2-VL-2B with deterministic decoding—evaluated via scalar accuracy, micro-averaged F1 for list attributes, and severity MAE/RMSE on a fixed validation split. We publicly release the annotation, and analysis scripts, including graph construction and evaluation scripts, to enable explainable, data-centric workflows for future intelligent vehicles.
@@ I Introduction
We introduce CAR-Scenes, a frame-level dataset designed for training and evaluating VLMs on interpretable scene understanding. CAR-Scenes consolidates 5,192 images from Argoverse1, Cityscapes, KITTI, and nuScenes under a 28-key category/sub-category knowledge base ( 350+ leaf attributes) that covers environment, road geometry, background-vehicle behavior, ego-vehicle behavior, vulnerable road users, sensor states, and a discrete severity scale (1–10). Labels are produced via a GPT-4o–assisted vision–language pipeline with human-in-the-loop verification; we release the exact prompts, post-processing rules, and per-field model performance for transparency and reuse.
@@ II Related Work
CAR-Scenes (our work) Single RGB frames Structured scene KB (28 keys; Severity) 5,192 frames 4,692/500 train/val Argoverse1, Cityscapes, KITTI, nuScenes Images + JSONL schema+validator+loader
@@
TABLE I: Related driving VLM datasets compared to CAR-Scenes.
Autonomous-driving datasets. KITTI, Cityscapes, BDD100K, Argoverse, nuScenes, Lyft Level 5, and Waymo Open provide multimodal data with object/pixel-aligned labels for detection, tracking, and mapping. These corpora include limited, coarse scene metadata and are not designed for schema-consistent scene-level semantics. CAR-Scenes complements them by releasing images + JSONL scene annotations for interpretable analysis rather than leaderboard evaluation.
Vision–language for driving. BDD-X, Talk2Car, DriveLM/DriveLM-Data, NuInstruct, and nuScenes-QA pair scenes with free-form text, QA, or graph-QA signals. While powerful, supervision and vocabularies are heterogeneous. CAR-Scenes instead offers a versioned, fixed schema (scene attributes incl. severity) enabling consistent querying and symbolic analysis.
Scene graphs & risk corpora. Visual Genome [RefWorks:krishna2017visual], RefCOCO [RefWorks:chen2025revisiting], and GQA [RefWorks:hudson2019gqa] probe general-image relations/referring or VQA [RefWorks:antol2015vqa], while Scenic [RefWorks:fremont2019scenic], RiskBench [RefWorks:kung2024riskbench], DReyeVE [RefWorks:alletto2016dr], and DeepScenario [RefWorks:lu2023deepscenario] center on simulators, trajectories, or attention—none provide a per-image, driving-specific semantic schema (signals, lane/lane-use, ego context) grounded directly in real images. As a result, they don’t support standardized, per-key evaluation of scene-level attributes or lightweight cross-dataset curation. CAR-Scenes addresses this by offering single-frame, schema-aligned annotations for environment, traffic control, road geometry, actors, and a coarse difficulty/severity proxy, with a consistent evaluation protocol that complements these efforts.
@@ III-A Sources and Scope
CAR-Scenes augments existing autonomous-driving datasets with scene-level annotations and a discrete severity score (1–10). We aggregate 5,192 images from Argoverse1, Cityscapes, KITTI, and nuScenes, chosen for complementary coverage of geography, weather, and urban form. We release annotations and tools only; original imagery remains under the licenses of the source datasets.
@@ III-B Schema and Knowledge Base
Design goals. CAR-Scenes targets scene-level, interpretable supervision that is (i) human-readable and stable across datasets, (ii) VLM-ready for training and evaluation, and (iii) schema-faithful with explicit enumerations and constraints.
@@ IV Dataset Statistics and Structure
This section quantifies the scale and semantic coverage of CAR-Scenes and documents how frames are sampled from source datasets. Our goal is to characterize attribute diversity and severity (1–10) distributions in a way that is directly useful for training and evaluating VLMs.
@@ IV-A Data Sources and Splits
Sources. CAR-Scenes comprises annotated frames from four widely used autonomous-driving datasets: Argoverse1, KITTI, Cityscapes, and nuScenes. These sources jointly cover varied geographies, road types, traffic mixes, and environmental conditions.
@@
CAR-Scenes includes 5,192 frames sampled from Argoverse1, Cityscapes, KITTI, and nuScenes. We provide a training split of 4,692 images and a fixed validation split of 500 images. Users may re-partition by source/city/condition using provided scripts.
@@ IV-G Licensing
CAR-Scenes combines our original materials (annotations, schema, prompt templates, validator, and code) with images from prior datasets (Cityscapes, KITTI, Argoverse v1, nuScenes). We license our annotations/schema/prompts under CC BY 4.0 and our code under MIT; these licenses apply only to our contributions. Images remain under their original licenses and are not re-licensed by us. We do not redistribute Cityscapes or nuScenes images; we provide annotations keyed to official filenames and require users to obtain those images from the original portals under their Terms of Use (nuScenes commercial use requires a separate license from Motional). KITTI and Argoverse v1 images may be redistributed only under their upstream CC BY-NC-SA licenses (KITTI: 3.0; Argoverse v1: 4.0) with attribution; if you include these images, the Non-Commercial + ShareAlike terms flow down to your bundle—otherwise release annotations-only. Consequently, any archive that includes KITTI/Argoverse imagery is non-commercial; for commercial use, exclude those images and pair our annotations with user-downloaded copies from the originals. Please cite the original datasets alongside this work.
@@ V Baseline Vision-Language Model
This section documents the Qwen2-VL-2B-Instruct baseline that accompanies the CAR-Scenes dataset release. Our goal is to provide an immediately reproducible reference model together with the exact training, validation, and evaluation protocol. An abstract of the per-field model performance is given in Table  LABEL:tab:llmvshuman.
@@ VI Conclusion
We introduced CAR-Scenes, a lightweight, image-grounded semantic dataset for vision–language modeling in autonomous driving that augments 5,192 frames from standard benchmarks with a 28-key schema (350+ attributes) and a discrete Severity label (1–10). Using a deterministic VLM pipeline with human verification, we release the schema, prompts, a validator, and a reproducible baseline intended to calibrate task difficulty (per-key accuracy/F1 and severity regression) rather than to set a leaderboard. CAR-Scenes enables 1) structured-field training/evaluation, compositional generalization over list-valued attributes (e.g., VRUs & night & flashing signal), 2) coverage auditing and curation across time-of-day, weather, and infrastructure, long-tail retrieval and triage via schema queries, cross-dataset transfer under a unified ontology, curriculum/active learning using severity as a difficulty signal, and 3) scenario/test selection by mapping schema constraints to downstream simulation or closed-loop evaluations—all without requiring video or simulators.

### v2

# CARScenes: Semantic VLM Dataset for Safe Autonomous Driving

@@ Abstract
CARScenes is a frame-level dataset for autonomous driving that enables the training and evaluation of vision–language models (VLMs) for interpretable, scene-level understanding. We annotate 5,192 images drawn from Argoverse1, Cityscapes, KITTI, and nuScenes using a 28-key category/sub-category knowledge base covering environment, road geometry, background-vehicle behavior, ego-vehicle behavior, vulnerable road users, sensor states, and a discrete severity scale (1–10), totaling 350+ leaf attributes. Labels are produced by a GPT-4o–assisted vision–language pipeline with human-in-the-loop verification; we release the exact prompts, post-processing rules, and per-field baseline model performance. CARScenes also provides attribute co-occurrence graphs and JSONL records that support semantic retrieval, dataset triage, and risk-aware scenario mining across sources. To calibrate task difficulty, we include reproducible, non-benchmark baselines—notably a LoRA-tuned Qwen2-VL-2B with deterministic decoding—evaluated via scalar accuracy, micro-averaged F1 for list attributes, and severity MAE/RMSE on a fixed validation split. We publicly release the annotation, and analysis scripts, including graph construction and evaluation scripts, to enable explainable, data-centric workflows for future intelligent vehicles.
@@ I Introduction
We introduce CARScenes, a frame-level dataset designed for training and evaluating VLMs on interpretable scene understanding. CARScenes consolidates 5,192 images from Argoverse1, Cityscapes, KITTI, and nuScenes under a 28-key category/sub-category knowledge base ( 350+ leaf attributes) that covers environment, road geometry, background-vehicle behavior, ego-vehicle behavior, vulnerable road users, sensor states, and a discrete severity scale (1–10). Labels are produced via a GPT-4o–assisted vision–language pipeline with human-in-the-loop verification; we release the exact prompts, post-processing rules, and per-field model performance for transparency and reuse.
@@ II Related Work
CARScenes (our work) Single RGB frames Structured scene KB (28 keys; Severity) 5,192 frames 4,692/500 train/val Argoverse1, Cityscapes, KITTI, nuScenes Images + JSONL schema+validator+loader
@@
TABLE I: Related driving VLM datasets compared to CARScenes.
Autonomous-driving datasets. KITTI, Cityscapes, BDD100K, Argoverse, nuScenes, Lyft Level 5, and Waymo Open provide multimodal data with object/pixel-aligned labels for detection, tracking, and mapping. These corpora include limited, coarse scene metadata and are not designed for schema-consistent scene-level semantics. CARScenes complements them by releasing images + JSONL scene annotations for interpretable analysis rather than leaderboard evaluation.
Vision–language for driving. BDD-X, Talk2Car, DriveLM/DriveLM-Data, NuInstruct, and nuScenes-QA pair scenes with free-form text, QA, or graph-QA signals. While powerful, supervision and vocabularies are heterogeneous. CARScenes instead offers a versioned, fixed schema (scene attributes incl. severity) enabling consistent querying and symbolic analysis.
Scene graphs & risk corpora. Visual Genome [RefWorks:krishna2017visual], RefCOCO [RefWorks:chen2025revisiting], and GQA [RefWorks:hudson2019gqa] probe general-image relations/referring or VQA [RefWorks:antol2015vqa], while Scenic [RefWorks:fremont2019scenic], RiskBench [RefWorks:kung2024riskbench], DReyeVE [RefWorks:alletto2016dr], and DeepScenario [RefWorks:lu2023deepscenario] center on simulators, trajectories, or attention—none provide a per-image, driving-specific semantic schema (signals, lane/lane-use, ego context) grounded directly in real images. As a result, they don’t support standardized, per-key evaluation of scene-level attributes or lightweight cross-dataset curation. CARScenes addresses this by offering single-frame, schema-aligned annotations for environment, traffic control, road geometry, actors, and a coarse difficulty/severity proxy, with a consistent evaluation protocol that complements these efforts.
@@ III-A Sources and Scope
CARScenes augments existing autonomous-driving datasets with scene-level annotations and a discrete severity score (1–10). We aggregate 5,192 images from Argoverse1, Cityscapes, KITTI, and nuScenes, chosen for complementary coverage of geography, weather, and urban form. We release annotations and tools only; original imagery remains under the licenses of the source datasets.
@@ III-B Schema and Knowledge Base
Design goals. CARScenes targets scene-level, interpretable supervision that is (i) human-readable and stable across datasets, (ii) VLM-ready for training and evaluation, and (iii) schema-faithful with explicit enumerations and constraints.
@@ IV Dataset Statistics and Structure
This section quantifies the scale and semantic coverage of CARScenes and documents how frames are sampled from source datasets. Our goal is to characterize attribute diversity and severity (1–10) distributions in a way that is directly useful for training and evaluating VLMs.
@@ IV-A Data Sources and Splits
Sources. CARScenes comprises annotated frames from four widely used autonomous-driving datasets: Argoverse1, KITTI, Cityscapes, and nuScenes. These sources jointly cover varied geographies, road types, traffic mixes, and environmental conditions.
@@
CARScenes includes 5,192 frames sampled from Argoverse1, Cityscapes, KITTI, and nuScenes. We provide a training split of 4,692 images and a fixed validation split of 500 images. Users may re-partition by source/city/condition using provided scripts.
@@ IV-G Licensing
CARScenes combines our original materials (annotations, schema, prompt templates, validator, and code) with images from prior datasets (Cityscapes, KITTI, Argoverse v1, nuScenes). We license our annotations/schema/prompts under CC BY 4.0 and our code under MIT; these licenses apply only to our contributions. Images remain under their original licenses and are not re-licensed by us. We do not redistribute Cityscapes or nuScenes images; we provide annotations keyed to official filenames and require users to obtain those images from the original portals under their Terms of Use (nuScenes commercial use requires a separate license from Motional). KITTI and Argoverse v1 images may be redistributed only under their upstream CC BY-NC-SA licenses (KITTI: 3.0; Argoverse v1: 4.0) with attribution; if you include these images, the Non-Commercial + ShareAlike terms flow down to your bundle—otherwise release annotations-only. Consequently, any archive that includes KITTI/Argoverse imagery is non-commercial; for commercial use, exclude those images and pair our annotations with user-downloaded copies from the originals. Please cite the original datasets alongside this work.
@@ V Baseline Vision-Language Model
This section documents the Qwen2-VL-2B-Instruct baseline that accompanies the CARScenes dataset release. Our goal is to provide an immediately reproducible reference model together with the exact training, validation, and evaluation protocol. An abstract of the per-field model performance is given in Table  LABEL:tab:llmvshuman.
@@ VI Conclusion
We introduced CARScenes, a lightweight, image-grounded semantic dataset for vision–language modeling in autonomous driving that augments 5,192 frames from standard benchmarks with a 28-key schema (350+ attributes) and a discrete Severity label (1–10). Using a deterministic VLM pipeline with human verification, we release the schema, prompts, a validator, and a reproducible baseline intended to calibrate task difficulty (per-key accuracy/F1 and severity regression) rather than to set a leaderboard. CARScenes enables 1) structured-field training/evaluation, compositional generalization over list-valued attributes (e.g., VRUs & night & flashing signal), 2) coverage auditing and curation across time-of-day, weather, and infrastructure, long-tail retrieval and triage via schema queries, cross-dataset transfer under a unified ontology, curriculum/active learning using severity as a difficulty signal, and 3) scenario/test selection by mapping schema constraints to downstream simulation or closed-loop evaluations—all without requiring video or simulators.